### v4

# Static quantification of player value for fantasy basketball

Abstract
Fantasy basketball has a rich underlying mathematical structure which makes optimal drafting strategy unclear. A central issue for category leagues is how to aggregate a player’s statistics from all categories into a single number representing general value. It is shown that under a simplified model of fantasy basketball, a novel metric dubbed the “G-score” is appropriate for this purpose. The traditional metric used by analysts, “Z-score”, is a special case of the G-score under the condition that future player performances are known exactly. The distinction between Z-score and G-score is particularly meaningful for head-to-head formats, because there is a large degree of uncertainty in player performance from one week to another. Simulated fantasy basketball seasons with head-to-head scoring provide evidence that G-scores do in fact outperform Z-scores in that context.
1 Introduction
1.1 What is fantasy basketball?
Fantasy basketball is a game in which participants manage virtual basketball teams based on the performances of real players. According to numbers from the Fantasy Sports & Gaming Association, approximately 20 million Americans and Canadians played fantasy basketball in 2022 [1]
1.1.1 Player selection
In fantasy basketball, “managers” draft actual players to their “teams” before the NBA or WNBA season. Like in a real sports draft, no player can be selected more than once, and once a player gets picked by a manager they are removed from the draft board. Usually, the picks are made in a snaking order as depicted in Table 1. Alternatively, players can be chosen via a mock auction for which all managers start with the same amount of virtual money [2].
Teams all have the same number of players, often thirteen. During the season, managers can swap their players with “free agents” or trade them to other managers. The exact rules for how managers are allowed to change their team during the season differ from league to league
Team A Team B Team C Team D Team E Team F
1 2 3 4 5 6
12 11 10 9 8 7
13 14 15 16 17 18
24 23 22 21 20 19
Table 1: Draft order for a six-team, four-round snake draft
1.1.2 Scoring
Fantasy basketball has several common scoring formats. As described by ESPN, the category-based formats are
  • Rotisserie: Rotisserie, or “Roto”, is the most common way to play fantasy basketball. In this scoring type, teams are ranked from first to last in each statistical category. Points are then awarded according to the order in each category and totaled to determine an overall score and league rank.
  • Head-to-Head: Each Category: H2H Each Category is the most common type of head-to-head play in fantasy basketball. For each scoring period, team totals are accumulated and a win, loss or tie is credited in each category based on the matchup results (i.e. 6-3-1 in a 10 category league) .
  • Head-to-Head: Most Categories: For each scoring period, team totals are accumulated in each of the categories. At the end of the scoring period, the winner is determined by which team wins the most number of categories. The end result is a win (1-0-0) , loss (0-1-0) or tie (0-0-1) . These results correspond directly to each team’s overall record [3].
Common settings specify nine categories: points, rebounds, assists, steals, blocks, three-pointers, field goal %, free throw %, and turnovers. Player performances are aggregated to get overall team metrics for each scoring period. Points, rebounds, assists, steals, blocks, three-pointers, and turnovers, henceforth dubbed “counting statistics” are aggregated by taking the sum. Field goal % and free throw %, henceforth dubbed “percentage statistics” are aggregated by taking the total number of successes over the total number of attempts for the category. For Rotisserie, player performances are aggregated over the course of an entire season. For head-to-head, scoring periods are usually weeks. Categories are won by the team with the higher metric, except for the turnovers category which is won by the team with the lower metric. Examples of how the scoring system works are included as Tables 2, 3, 4(a), and 4(b)
Points Turnovers Free throw %
Team A Player 1 10 1 66% (2/3)
Player 2 15 2 100% (4/4)
Player 3 33 2 50% (5/10)
Aggregate 58 5 64% (11/17)
Team B Player 4 12 1 50% (1/2)
Player 5 8 1 0% (0/1)
Player 6 20 1 80% (8/10)
Aggregate 40 3 69% (9/14)
Table 2: Example of how metrics are aggregated and compared. Team A wins points and loses the two other categories. It loses turnovers because for that category, the team with the lower total wins
Points rank Blocks rank Rebounds rank Total Result
Team A 1 3 3 7 Second
Team B 2 2 1 5 First
Team C 3 1 4 8 Third
Team D 4 4 2 10 Fourth
Table 3: Example of rotisserie scoring with four teams and three categories. Ranks are calculated by totals across a full season- 1 is best, 4 is worst
Week 1 Week 2 Week 3 Total Result
Team A 5-4 1-8 4-5 10-17 Fourth
Team B 4-5 4-5 3-6 11-16 Third
Team C 3-6 8-1 6-3 17-10 First
Team D 6-3 5-4 5-4 16-11 Second
(a) Each Category scoring
Week 1 Week 2 Week 3 Total Result
Team A 5-4 (W) 1-8 (L) 4-5 (L) 1-2 Third
Team B 4-5 (L) 4-5 (L) 3-6 (L) 0-3 Fourth
Team C 3-6 (L) 8-1 (W) 6-3 (W) 2-1 Second
Team D 6-3 (W) 5-4 (W) 5-4 (W) 3-0 First
(b) Most Category scoring
Table 4: Scoring with four teams, nine categories, and three weeks. Teams face each other in a round robin order
1.1.3 Position rules
On a daily basis, players must be placed at positions in order for their games to count. In fantasy basketball the main positions are center, point guard, shooting guard, power forward, and small forward. Each player has a set of positions for which they are eligible. Teams are limited to a certain number of players at each position for day, and when no slots are available, players must be placed “on the bench” where their games do not count
1.2 Static value assessment
The fantasy basketball problem is often reduced to a matter of static player values independent of context. Many fantasy basketball experts e.g. Basketball Monster publish general value estimates before each season, based on how they expect each player to perform [4]. Managers can then use these to decide which players to pick during their snake drafts. Managers can also use static value estimates for auctions, converting them to dollar values by subtracting out replacement value (the value of a player that can be added for free after the auction) then re-scaling to the amount of money available [5].
Quantifying value for category leagues is challenging because there is no obvious way to combine performances across different categories. The typical method used by Basketball Monster and others is to calculate Z-scores of expected performances by subtracting the mean for the category and dividing by standard deviation, then summing the categorical Z-scores for each player [6].
As will be shown for a simplified model of fantasy basketball, this “Z-score” method can be justified, but only as a special case of the more generally applicable “G-score” method. For the model of fantasy basketball that will be presented, G-scores are a more complete solution
2 Justifying G-scores
2.1 Model
Consider a manager whose turn it is to select a player. Assume that
  • Each managers’ objective $V$ is to have the highest possible expected value of number of categories won against an arbitrary opponent, across an arbitrary scoring period. The set of categories is $C$, of size $|C|$
  • All players on all teams, except for the one being currently selected, are chosen randomly from set $Q$. $Q$ is composed of $|Q|$ players, where $|Q|$ is the number of players that will be drafted into the league. E.g. if twelve managers are each picking thirteen players, $|Q|=12*13=156$
  • Performances for all players are chosen randomly from a set of possible outcomes $W$ of size $|W|$. The selection process is independent for each player
  • All games played by all chosen players, $N$ per team, count towards team totals
  • A player’s value is proportional to $V_{p}$ $V_{0}$, where $V_{p}$ is the objective value given that player $p$ is chosen, and $V_{0}=\frac{|C|}{2}$, the default objective value
The question is how player value should be calculated under these conditions. How well the conditions represent real fantasy basketball is discussed in Section 3.1
2.2 Solution
It is posited that under the conditions of the simplified model of fantasy basketball laid out in Section 2.1, a novel system dubbed G-score is appropriate for quantifying player value. G-scores per category are defined by Figure 5(c), and their sum is the aggregate G-score for a player. Essentially, G-scores are the same as Z-scores except the denominators are modified to add period-to-period variance to the player-to-player-variance already included in Z-scores. Z-scores are equivalent to G-scores with $|W|=1$, removing all period-to-period variance.
$\theta_{M}\left(q,w\right)$ Performance for player $q$ in $w$
$\mu_{M}\left(q\right)$ Mean of $\theta_{M}\left(q,w\right)$ for player $q$ across $W$: $\frac{\sum_{w\in W}\theta_{M}\left(q,w\right)}{|W|}$
$\mu_{M}$ Mean of $\mu_{M}\left(q\right)$ across players: $\frac{\sum_{q\in Q}\mu_{M}\left(q\right)}{|Q|}$
$\sigma_{M}$ Standard deviation of $\mu_{M}\left(q\right)$ across $Q$: $\sqrt{\frac{\sum_{q\in Q}\left(\mu_{M}\left(q\right)-\mu_{M}\right)^{2}}{|Q|}}$
$\tau_{M}\left(q\right)$ Standard deviation of $\theta_{M}\left(q,w\right)$ across $W$: $\sqrt{\frac{\sum_{w\in W}\left(\theta_{M}\left(q,w\right)-\mu_{M}\left(q\right)\right)^{2}}{|W|}}$
$\tau_{M}$ Root-mean-square of $\tau_{M}\left(q\right)$: $\sqrt{\frac{\sum_{q\in Q}\tau_{M}\left(q\right)^{2}}{|Q|}}$
$\kappa$ Kappa factor: $\frac{2N}{2N-1}$
(a) Definitions for counting statistics, used for G-scores
$\theta_{A}\left(q,w\right)$ Number of attempts made by player $q$ in $w$
$\theta_{R}\left(q,w\right)$ Success rate for player $q$ in $w$
$\mu_{A}\left(q\right)$ Mean of $\theta_{A}\left(q,w\right)$ for player $q$ across $W$: $\frac{\sum_{w\in W}\theta_{A}\left(q,w\right)}{|W|}$
$\mu_{A}$ Mean of $\mu_{A}\left(q\right)$ across players: $\frac{\sum_{q\in Q}\mu_{A}\left(q\right)}{|Q|}$
$\mu_{R}\left(q\right)$ Composite success rate for player $q$ across $W$: $\frac{\sum_{w\in W}\theta_{A}\left(q,w\right)\theta_{R}\left(q,w\right)}{\sum_{w\in W}\theta_{A}\left(q,w\right)}$
$\mu_{R}$ Composite success rate across players: $\frac{\sum_{q\in Q}a_{q}\mu_{R}\left(q\right)}{\sum_{q\in Q}a_{q}}$
$\sigma_{R}$ Standard deviation of $\mu_{R}\left(q\right)$ across $Q$: $\sqrt{\frac{\sum_{q\in Q}\left(\mu_{R}\left(q\right)-\mu_{R}\right)^{2}}{|Q|}}$
$\tau_{R}\left(q\right)$ Standard deviation of $\frac{\theta_{A}\left(q,w\right)}{\mu_{A}}*\left(\theta_{R}\left(q,w\right)-\mu_{R}\right)$ across $W$ for player q
$\tau_{R}$ Root-mean-square of $\tau_{R}\left(q\right)$: $\sqrt{\frac{\sum_{q\in Q}\tau_{R}\left(q\right)^{2}}{|Q|}}$
(b) Definitions for percentage statistics, used for G-scores
Points/Rebounds/Assists/Steals/Blocks/Threes Turnovers FG % / FT %
$\frac{\mu_{M}\left(q\right)-\mu_{M}}{\sqrt{\left(\sigma_{M}^{2}+\kappa*\tau_{M}^{2}\right)}}$ $\frac{\mu_{M}-\mu_{M}\left(q\right)}{\sqrt{\left(\sigma_{M}^{2}+\kappa*\tau_{M}^{2}\right)}}$ $\frac{\frac{a_{q}}{\mu_{A}}*\left(\mu_{R}\left(q\right)-\mu_{R}\right)}{\sqrt{r_{\sigma}^{2}+\kappa*\tau_{R}^{2}}}$
(c) Definition of G-score per category. The aggregate G-score is the sum across categories
Table 5: Definitions for G-scoring
A detailed proof for why the G-score solves the problem is included in Appendix A. More briefly, all players contribute period-to-period variance $\tau_{M}^{2}$, and all players except the player in question contribute player-to-player variance $\sigma_{M}^{2}$. Invoking the central limit theorem, the point differential between two teams for a counting statistic can then be written as
$\mathcal{N}\left(\mu_{M}-\mu_{M}\left(p\right),\left(2N-1\right)*\sigma_{M}^{2}+\left(2N\right)*\tau_{M}^{2}\right)$
If this value is below zero, than team $A$ scores more points than team $B$ and therefore wins the category. The probability that this value is below zero is the cumulative distribution function of the normal distribution at zero. Therefore, the probability of victory is the cumulative distribution function of the distribution at zero, which is approximately
$\frac{1}{2}\left[1+\frac{1}{\sqrt{\pi*\left(N-\frac{1}{2}\right)}}*\frac{\mu_{M}\left(p\right)-\mu_{M}}{\sqrt{\sigma_{M}^{2}+\frac{2N}{2N-1}*\tau_{M}^{2}}}\right]$
This includes the formula for G-score. Applying the same procedure to the percentage statistics is more complicated, but has an equivalent result. The manager’s objective function $V$ is the sum of winning probabilities across categories, so it is
$V=\frac{1}{2}\left[|C|+\frac{1}{\sqrt{\pi*\left(N-\frac{1}{2}\right)}}*\sum_{c\in C}G_{pc}\right]$
Where $G_{pc}$ is player $p$’s G-score for category $c$. Subtracting out the default expected number of wins yields
$V-V^{\prime}=\frac{1}{2\sqrt{\pi*\left(N-\frac{1}{2}\right)}}*\sum_{c\in C}G_{pc}$
This is player value, as defined in Section 2.1. G-score is clearly proportional to it, so G-score meets the criteria for player value as well
2.3 Simulation
Simulations were run to test the efficacy of G-scoring. The simulated seasons were twelve-team, thirteen-player head-to-head competitions. Each player’s performance in each simulated season was generated by randomly sampling twenty weeks from their actual performances, excluding weeks for which they were injured. Players were only included if they had at least 10 non-injured weeks. Teams were paired against each other via round robin format, as they are in real head to head competition, and winners were decided by which team had the most points by the end of the twenty week season. All games played by all players counted. Managers had access to weekly performance numbers for each player, allowing them to calculate all relevant metrics for all players. Z-scores based on the full league were used to choose $Q$, based on which managers calculated Z-scores and G-scores. The $\kappa$ factor was approximated at $1$. For the Z vs G matchup, Z-scoring was was tried at each “seat” (position in the draft order) of a snake draft, against a field of drafters using G-score. One thousand of the twenty-week simulated seasons were run for every draft seat, allowing for robust estimates of how well that strategy really would have performed in that situation with error bars no greater than
$\frac{\sqrt{1000*\frac{1}{2}{(1-\frac{1}{2})}}}{1000}\approx 1.6\%$
The same was also done in reverse, with G-score as the single drafter against a field of Z-score drafters.
Z vs. G G vs. Z
Most Categories Each Category Most Categories Each Category
Aggregate 0.4% $0.5\%$ $32.5\%$ $21.4\%$
Seat 0 $0.5\%$ $0.8\%$ $41.6\%$ $30.7\%$
Seat 1 $0.2\%$ $0.3\%$ $21.1\%$ $14.5\%$
Seat 2 $0.3\%$ $0.8\%$ $33.7\%$ $29.2\%$
Seat 3 $0.5\%$ $0.7\%$ $30.7\%$ $19.1\%$
Seat 4 $0.3\%$ $0.4\%$ $22.7\%$ $14.1\%$
Seat 5 $0.0\%$ $0.1\%$ $25.6\%$ $14.5\%$
Seat 6 $0.1\%$ $0.4\%$ $30.6\%$ $17.2\%$
Seat 7 $0.4\%$ $0.1\%$ $53.0\%$ $33.3\%$
Seat 8 $0.2\%$ $0.0\%$ $48.8\%$ $38.6\%$
Seat 9 $0.5\%$ $0.0\%$ $33.4\%$ $16.9\%$
Seat 10 $1.1\%$ $0.8\%$ $22.2\%$ $11.4\%$
Seat 11 $0.8\%$ $1.4\%$ $26.8\%$ $17.1\%$
Table 6: Win rates of Z-scores and G-scores against each other
3 Discussion
3.1 Model assumptions
Of course, the model presented in Section 2.1 to justify G-scores is not a perfect replica of how fantasy basketball works in practice. However, it does capture the main considerations, and is more feasible to solve than alternatives. This justifies it as a model worthy of study
3.1.1 Examining model assumptions
While no manager would pick players truly at random, if one picked players purely based on a ranking list with no consideration for team fit, their per-category totals would be functionally mostly random. This makes fully random selection a decent proxy for other drafters picking via ranking list, which would be much more complicated to analyze as discussed in LABEL:Comp.
The set of possible performances $W$ represents a manager’s belief about the probability distributions underlying future performance. $|W|$ can be infinite, making the probability distribution continuous, so any belief about player performance can be encoded with this model. The only sense in which this distorts real fantasy basketball is that it does not account for performance distribution changes throughout a season. See Appendix C.1 and C.2 for discussion on how to handle that complexity.
Ignoring position requirements is also not a terrible assumption for fantasy basketball, since position requirements are usually loose. See Appendix B for more discussion.
Managers generally want to maximize the expected value of their number of categories won. That is somewhat skewed from the real objective for Most Categories, but with a static ranking list there is no way of incorporating which categories a team is already weak or strong in, so maximizing for expected value is a reasonable proxy.
Finally, the idea of a value system is to use it to determine whether a player or group of players has more value than another. Ideally this calculation would consider player values in the context of each other, but that is impossible with a static system. Therefore, the only sensible way to aggregate value is to add marginal impacts of individual players
3.1.2 Computational complexity of game theory solution
Theoretically, fantasy basketball can be modeled as a game wherein managers choose their own ranking lists and draft players accordingly. That game mimics real fantasy basketball better than the model laid out in Section 2.1, since it obviates the need to assume that other players are chosen randomly. It is also a finite game, so by Nash’s theorem it has a Nash Equilibrium (an equilibrium from which no manager has incentive to deviate). Calculating this Nash Equilibrium would be a more precise way to solve the static value quantification problem. However, performing the calculation would be computationally intractable.
To calculate the Nash equilibrium, it would be necessary to know how each manager would fare under each possible strategy profile. Say that there were twelve teams with thirteen picks each, and there were $500$ total candidates. The number of possible rankings of relevant players would be
$\frac{500!}{\left(500-156\right)!}\approx 10^{409}$
No modern computer is capable of this many operations in any amount of time. So while theoretically a Nash Equilibrium would exist, it would not be possible to actually calculate it
3.2 Consequences for Head-to-head and Rotisserie
Even assuming that long-term means are known perfectly, performances are volatile from from one week to another. This leads $\tau_{M}^{2}$ to be large for head to head formats, re-weighting G-scores in a significant way relative to Z-scores. Steals in particular are down-weighted heavily as demonstrated by Table 7 based on real data from the 2022-2023 season.
$\sigma_{M}^{2}$ $\tau_{M}^{2}$ Z-score denominator G-score denominator G-score as fraction of Z-score
Assists 41.57 31.55 6.45 8.55 75%
Blocks 2.35 2.70 1.53 2.25 68%
Field Goal % 0.003 0.007 0.06 0.10 56%
Free Throw % 0.009 0.018 0.10 0.16 58%
Points 325.52 448.32 18.04 27.82 65%
Rebounds 52.01 57.55 7.21 10.47 69%
Steals 1.01 4.20 1.00 2.28 44%
Threes 9.52 9.04 3.09 4.31 72%
Turnovers 5.45 8.85 2.34 3.78 62%
Table 7: Empirical values for denominators of Z-score and G-score with real data from the 2022-2023 season, assuming random sampling from all weeks of the NBA season. $Q$ was calculated as the top 156 players by base Z-score, determined across the entire league
Z-scores are not as problematic for Rotisserie, because weekly variations do not matter across a full season. The result is that $\tau_{M}^{2}$ is likely small relative to $\sigma_{M}^{2}$, so dropping it as Z-scores do is not as problematic. Still, using Z-scores for Rotisserie is not precisely optimal, since they ignore that some categories are harder to forecast for a full season than others
3.3 Intuition
Implicitly, Z-scores are accounting for the fact that other players are unknowns, and therefore they introduce variance in performance means. Z-scores downweight the categories that have the most uncertainty across players, and upweight the categories that are most consistent.
This logic is reasonable but also incomplete, in the sense that if player selections are unknowns, then player performances should be unknowns as well. G-scores are a version of Z-scores fixed to account for this additional uncertainty. They downweight categories that are volatile in terms of performances in the same way that Z-scores downweight categories that are volatile across players
For intuition on why this re-weighting makes sense, consider a category which has significant uncertainty across players. A Z-score advantage in that category is worth relatively little, because random chance in other player choices has a relatively large effect. There is more marginal benefit to investing in a more stable category, for which a slight advantage can translate into consistent wins. G-scoring accounts for this and makes it so that an additional point of value is worth the same for each category. When all managers are using it, it creates a sort of equilibrium, where no manager can gain an advantage by targeting less volatile categories
3.4 General utility of static ranking lists
While convenient, static ranking lists are also fundamentally suboptimal, because they preclude the possibility of adapting based on draft circumstances.
One common adaptive strategy is called punting. When a manager punts, they strategically sacrifice one category in order to gain a consistent edge in others. To punt effectively, a manager needs to adapt to the players that happen to be available and form a punting strategy around them. For example, if a player like Giannis Antetokounmpo who is generally strong but weak in free throw percentage is available, it suggests that punting free throws could be a profitable strategy. The same strategy would not work nearly as well if a player like Giannis was not available. It cannot be known beforehand whether Giannis will be available to draft, making it impossible to craft an optimal punting strategy before the draft has begun. Static ranking lists therefore cannot punt optimally.
When interpreting these results, it should be remembered both that the simulated model is incomplete, and that static ranking lists are fundamentally suboptimal. In a real draft, with managers making more optimal decisions, these results will not necessarily be applicable. Still, these simulations provide some reassurance that the basis of G-scores is sound
3.5 Simulation results
Table 6 demonstrates that drafters selecting players via Z-score consistently performed poorly against opponents using G-score, always scoring far below the baseline expected rate of $\frac{1}{12}$ or 8.33%. In the reverse scenario, G-score drafters performed much better. They won $21.4\%$ of simulated Each Category seasons, and $32.5\%$ of Most Categories seasons.
When interpreting these results, it should be remembered both that the simulated model is incomplete, and that static ranking lists are fundamentally suboptimal. In a real draft, with managers making more optimal decisions, these results will not necessarily be applicable
3.6 Applying of the G-score formula
3.6.1 Defining Q
There are a few sensible ways to define $Q$. One is to find a pseudo-equilibrium state in which $Q$ is composed of the top $|Q|$ players by G-score value. This can be accomplished by repeatedly re-calculating the top $|Q|$ players by G-score, setting $Q$ to them, and adjusting G-scores accordingly. If $Q$ does not change after an iteration, then it is an equilibrium set.
For simplicity, it is also possible to estimate $Q$ using a popular ranking metric
3.6.2 The $\kappa$ factor
The $\kappa$ factor accounts for the fact that the player currently being chosen contributes additional period-to-period variance. This necessitates the inclusion of a small $\frac{2N}{2N-1}$ adjustment factor to period-to-period variance.
Most leagues have twelve or thirteen players per team, so the $\kappa$ factor is generally between $\frac{26}{25}=1.040$ and $\frac{24}{23}\approx 1.043$. As a multiplier it is therefore quite small, and can be held constant at $1.04$ or removed entirely for the sake of simplicity
4 Conclusion
G-scores, an extension of Z-scores, are justified via a set of simple assumptions. This improvement is especially important for head-to-head formats, because weekly variance can be significant, and ignoring it is heavily sub-optimal.
There are many areas for future work, including optimizing picks dynamically based on drafting context. Static ranking lists are fundamentally suboptimal relative to dynamic strategies so a truly sophisticated treatment of fantasy basketball would model some level of adaptation.
Disclaimer: The views and opinions expressed in this article are those of the independent author and do not represent those of any organization, company or entity
Appendix A Detailed justification of G-score
A.1 Mathematical assumptions
Several assumptions are invoked to simplify the justification behind G-scores. They are
  • Category totals for a team can be estimated as Normal distributions due to the central limit theorem
  • The error function applied to a category total can be approximated with its first-order Taylor series
  • $\frac{X}{Y}$ can be approximated as a Normal distribution with $E\left(\frac{X}{Y}\right)\approx\frac{E(X)}{E(Y)}$. This assumes that $X$ and $Y$ are approximately Normal. independent, positive, and have low coefficient of variations [7]
  • The variance of a ratio of success over attempts can be estimated as the first three terms in its Taylor series
  • The period-to-period variance contribution of a player is approximately equal to the aggregate contribution
    $\tau_{M}\left(p\right)\approx\tau_{M}$ (1)
    $\tau_{R}\left(p\right)\approx\tau_{R}$ (2)
  • Two quantities are approximately independent. Specifically,
    $\frac{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)}{{\sum_{\left(q,w\right)\in P}\left(\theta_{A}\left(q,w\right)\right)+\theta_{A}\left(p,k\right)}}\perp\frac{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)\theta_{R}\left(q,w\right)-\sum_{q\in T_{A}}\theta_{A}\left(q,w\right)\mu_{R}}{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)}$ (3)
  • The expected number of free throw attempts by a team where one player is known is approximately equal to the number of free throw attempts by an arbitrary team. Mathematically,
    $\mu_{A}*\left(N-1\right)+\mu_{A}\left(p\right)\approx N\mu_{A}$ (4)
  • When calculating variance, the expected total number of free throw attempts by a team where one player is known is approximately equal to the number of free throws made by an arbitrary team. Mathematically,
    $\left(N-1\right)\mu_{A}\mu_{R}+\mu_{A}\left(p\right)\mu_{R}\left(p\right)\approx N\mu_{R}\mu_{A}$ (5)
A.2 Counting statistics
A weekly total for a random player is essentially a random draw from a two-dimensional set of player and weekly performance.
The mean of this draw is $\mu_{M}$. The variance can be calculated as
$\operatorname{Var}_{q,w}\left(\theta_{M}\left(q,w\right)\right)=\operatorname{E}_{q,w}\left(\left(\theta_{M}\left(q,w\right)-\mu_{M}\right)^{2}\right)=\operatorname{E}_{q}\left(\operatorname{E}_{w}\left(\left(\theta_{M}\left(q,w\right)-\mu_{M}\right)^{2}\right)\right)$
Since $\operatorname{E}\left(X^{2}\right)=\operatorname{Var}\left(X\right)+\operatorname{E}\left(X\right)^{2}$, this translates to
$\operatorname{Var}_{q,w}\left(\theta_{M}\left(q,w\right)\right)=\operatorname{E}_{q}\left(\operatorname{Var}_{w}\left(\theta_{M}\left(q,w\right)-\mu_{M}\right)+\left(\operatorname{E}_{w}\left(\theta_{M}\left(q,w\right)-\mu_{M}\right)\right)^{2}\right)$
$=\operatorname{E}_{q}\left(\operatorname{Var}_{w}\left(\theta_{M}\left(q,w\right)\right)+\left(\mu_{M}\left(q\right)-\mu_{M}\right)^{2}\right)$
$=\operatorname{E}_{q}\left(\tau_{M}\left(q\right)\right)+\operatorname{Var}_{q}\left(\mu_{M}\left(q\right)\right)+\operatorname{E}_{q}\left(\mu_{M}\left(q\right)-\mu_{M}\right)^{2}$
$=\operatorname{E}_{q}\left(\tau_{M}\left(q\right)\right)+\operatorname{Var}_{q}\left(\mu_{M}\left(q\right)\right)$
These two terms match the definitions of $\sigma_{M}^{2}$ and $\tau_{M}^{2}$ as defined by 5(a), leading to a variance of
$\tau_{M}^{2}+\sigma_{M}^{2}$
Team $B$’s $N$ players in aggregate are then distributed with
  • A mean of $N*\mu_{M}$
  • A variance of $N*\left(\sigma_{M}^{2}+\tau_{M}^{2}\right)$
Keeping in mind that the unchosen player’s performance has mean $\mu_{M}\left(p\right)$ and variance $\tau_{M}\left(p\right)^{2}$ by definition, team $A$’s distribution has
  • A mean of $\left(N-1\right)*\mu_{M}+\mu_{M}\left(p\right)$
  • A variance of $\left(N-1\right)*\left(\sigma_{M}^{2}+\tau_{M}^{2}\right)+\tau_{M}\left(p\right)^{2}$
Invoking the central limit theorem for team $B$’s total minus team $A$’s yields
$D\sim\mathcal{N}\left(\mu_{M}-\mu_{M}\left(p\right),\left(2N-1\right)*\left(\sigma_{M}^{2}+\tau_{M}^{2}\right)+\tau_{M}\left(p\right)^{2}\right)$
The CDF of this quantity at zero is the probability of team $A$ winning the category. It is
$\frac{1}{2}\left[1+\operatorname{erf}\left(\frac{\mu_{M}\left(p\right)-\mu_{M}}{\sqrt{2\left(\left(2N-1\right)*\left(\sigma_{M}^{2}+\tau_{M}^{2}\right)+\tau_{M}\left(p\right)^{2}\right)}}\right)\right]$
The first-order Taylor approximation, which is $\frac{2*X}{\sqrt{\pi}}$, can be used for the error function. This simplifies the expression to
$\frac{1}{2}\left[1+\frac{2}{\sqrt{\pi}}*\frac{\mu_{M}\left(p\right)-\mu_{M}}{\sqrt{2\left((2N-1)(\sigma_{M}^{2}+\tau_{M}^{2})+\tau_{M}\left(p\right)^{2}\right)}}\right]$
Using the approximation of equation 1 that $\tau_{M}\left(p\right)\approx\tau_{M}$ this becomes
$\frac{1}{2}\left[1+\frac{2}{\sqrt{\pi}}*\frac{\mu_{M}\left(p\right)-\mu_{M}}{\sqrt{2\left((2N-1)*\sigma_{M}^{2}+2N*\tau_{M}^{2}\right)}}\right]$
$\frac{1}{2}\left[1+\frac{2}{\sqrt{\pi}}*\frac{\mu_{M}\left(p\right)-\mu_{M}}{\sqrt{2\left(2N-1)\right)\left(\sigma_{M}^{2}+\frac{2N}{2N-1}*\tau_{M}^{2}\right)}}\right]$
$=\frac{1}{2}\left[1+\frac{1}{\sqrt{\pi*\left(N-\frac{1}{2}\right)}}*\frac{\mu_{M}\left(p\right)-\mu_{M}}{\sqrt{\sigma_{M}^{2}+\frac{2N}{2N-1}\tau_{M}^{2}}}\right]$
A.3 Percentage statistics
The percentage statistics are more complicated to model
A.3.1 Mean of the differential
Consider $P$ to be a set of pairs $\left(q,w\right)$ where $q$ is the chosen player and $w$ is the chosen week. Team $B$’s aggregate performance is a random draw from the two-dimensional distribution of player and scoring period
$O_{B}=\frac{\sum_{\left(q,w\right)\in P}\theta_{R}\left(q,w\right)\theta_{A}\left(q,w\right)}{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)}$
Each element of the sum is equivalent. So
$O_{\mu_{B}}=\frac{N*\operatorname{E}\left(\theta_{R}\left(q,w\right)\theta_{A}\left(q,w\right)\right)}{N*\operatorname{E}\left(\theta_{A}\left(q,w\right)\right)}=\frac{\operatorname{E}\left(\theta_{R}\left(q,w\right)\theta_{A}\left(q,w\right)\right)}{\operatorname{E}\left(\theta_{A}\left(q,w\right)\right)}=\frac{\operatorname{E}\left(\sum_{q\in Q,w\in W}\theta_{R}\left(q,w\right)\theta_{A}\left(q,w\right)\right)}{\operatorname{E}\left(\sum_{q\in Q,w\in W}\theta_{A}\left(q,w\right)\right)}$
Since $\mu_{A}\left(q\right)$ is defined as $\frac{\sum_{w\in W}\theta_{A}\left(q,w\right)}{|W|}$, this can be rewritten to
$O_{\mu_{B}}=\frac{\operatorname{E}\left(\sum_{q\in Q,w\in W}\theta_{A}\left(q,w\right)*\theta_{R}\left(q,w\right)\right)}{\operatorname{E}\left(|W|*\sum_{q\in Q}\mu_{A}\left(q\right)\right)}$
$\mu_{R}\left(q\right)$ is defined as $\frac{\sum_{w\in W}\theta_{R}\left(q,w\right)*\theta_{A}\left(q,w\right)}{\sum_{w}\theta_{A}\left(q,w\right)}$ or $\frac{\sum_{w\in W}\theta_{R}\left(q,w\right)*\theta_{A}\left(q,w\right)}{|W|*\mu_{A}\left(q\right)}$. That means $\sum_{w\in W}\theta_{R}\left(q,w\right)*\theta_{A}\left(q,w\right)=|W|*\mu_{A}\left(q\right)*\mu_{R}\left(q\right)$, allowing for a substitution to
$O_{\mu_{B}}=\frac{\operatorname{E}\left(\sum_{q\in Q}|W|*\mu_{A}\left(q\right)*\mu_{R}\left(q\right)\right)}{\operatorname{E}\left(|W|*\sum_{q\in Q}\mu_{A}\left(q\right)\right)}=\frac{\operatorname{E}\left(\sum_{q\in Q}\mu_{A}\left(q\right)*\mu_{R}\left(q\right)\right)}{\operatorname{E}\left(\sum_{q\in Q}\mu_{A}\left(q\right)\right)}=\mu_{R}$
Consider $k$ to be the randomly sampled week for the unchosen player $p$’s performance. Then, team A’s success rate $O_{A}$ is
$O_{A}=\frac{\sum_{\left(q,w\right)\in P}\theta_{R}\left(q,w\right)\theta_{A}\left(q,w\right)+\theta_{R}\left(k,p\right)\theta_{A}\left(p,k\right)}{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)+\theta_{A}\left(p,k\right)}$
$O_{A}$ can be equivalently rewritten as
$\frac{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)\left(\theta_{R}\left(q,w\right)-\mu_{R}\right)+\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)\mu_{R}+\theta_{A}\left(p,k\right)\left(\theta_{R}\left(k,p\right)-\mu_{R}\right)+\theta_{A}\left(p,k\right)\mu_{R}}{\sum_{\left(q,w\right)\in P}\mu_{A}\left(q\right)+\theta_{A}\left(p,k\right)}$
Or broken into separate terms
$O_{A}=$ $\frac{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)\mu_{R}+\theta_{A}\left(p,k\right)\mu_{R}}{{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)+\theta_{A}\left(p,k\right)}}+\frac{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)\left(\theta_{R}\left(q,w\right)-\mu_{R}\right)}{{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)+\theta_{A}\left(p,k\right)}}$
$+\frac{\theta_{A}\left(p,k\right)\left(\theta_{R}\left(k,p\right)-\mu_{R}\right)}{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)+\theta_{A}\left(p,k\right)}$
The first term can be simplified easily
$\mu_{R}\frac{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)+\theta_{A}\left(p,k\right)}{{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)+\theta_{A}\left(p,k\right)}}+\frac{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)\left(\theta_{R}\left(q,w\right)-\mu_{R}\right)}{{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)+\theta_{A}\left(p,k\right)}}$
$+\frac{\theta_{A}\left(p,k\right)\left(\theta_{R}\left(k,p\right)-\mu_{R}\right)}{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)+\theta_{A}\left(p,k\right)}$
$=$ $\mu_{R}+\frac{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)\left(\theta_{R}\left(q,w\right)-\mu_{R}\right)}{{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)+\theta_{A}\left(p,k\right)}}+\frac{\theta_{A}\left(p,k\right)\left(\theta_{R}\left(k,p\right)-\mu_{R}\right)}{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)+\theta_{A}\left(p,k\right)}$
The expected value of team $A$’s success rate is then
$O_{\mu_{A}}=\mu_{R}+\operatorname{E}\left(-\frac{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)\left(\theta_{R}\left(q,w\right)-\mu_{R}\right)}{{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)+\theta_{A}\left(p,k\right)}}-\frac{\theta_{A}\left(p,k\right)\left(\theta_{R}\left(k,p\right)-\mu_{R}\right)}{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)+\theta_{A}\left(p,k\right)}\right)$
To get the expected value of the differential, it is subtracted from $\mu_{R}$, leading to
$D_{\mu}=$ $-\operatorname{E}\left(\frac{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)\theta_{R}\left(q,w\right)-\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)\mu_{R}}{{\sum_{\left(q,w\right)\in P}\left(\theta_{A}\left(q,w\right)\right)+\theta_{A}\left(p,k\right)}}\right)$
$-\operatorname{E}\left(\frac{\theta_{A}\left(p,k\right)\left(\theta_{R}\left(p,k\right)-\mu_{R}\right)}{\sum_{\left(q,w\right)\in P}\left(\theta_{A}\left(q,w\right)\right)+\theta_{A}\left(p,k\right)}\right)$
$=$ $-\operatorname{E}\left(\frac{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)}{{\sum_{\left(q,w\right)\in P}\left(\theta_{A}\left(q,w\right)\right)+\theta_{A}\left(p,k\right)}}*\frac{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)\theta_{R}\left(q,w\right)-\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)\mu_{R}}{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)}\right)$
$-\operatorname{E}\left(\frac{\theta_{A}\left(p,k\right)\left(\theta_{R}\left(p,k\right)-\mu_{R}\right)}{\sum_{\left(q,w\right)\in P}\left(\theta_{A}\left(q,w\right)\right)+\theta_{A}\left(p,k\right)}\right)$
With the assumption of equation 3, this becomes
$D_{\mu}=$ $-\operatorname{E}\left(\frac{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)}{{\sum_{\left(q,w\right)\in P}\left(\theta_{A}\left(q,w\right)\right)+\theta_{A}\left(p,k\right)}}\right)$
$\quad*\operatorname{E}\left(\frac{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)\theta_{R}\left(q,w\right)-\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)\mu_{R}}{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)}\right)$
$-\operatorname{E}\left[\frac{\theta_{A}\left(p,k\right)\left(\theta_{R}\left(p,k\right)-\mu_{R}\right)}{\sum_{\left(q,w\right)\in P}\left(\theta_{A}\left(q,w\right)\right)+\theta_{A}\left(p,k\right)}\right]$
The expected value of $\frac{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)\theta_{R}\left(q,w\right)}{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)}$ is $\mu_{R}$ as proven for $O_{B}$, and $\frac{\sum_{q\in P}\theta_{A}\left(q,w\right)\mu_{R}}{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)}$ is $\mu_{R}$ because it can be pulled out of the sum as a constant. Therefore the first term cancels out and disappears. All that remains is the original second term,
$D_{\mu}=-\operatorname{E}\left(\frac{\theta_{A}\left(p,k\right)\left(\theta_{R}\left(k,p\right)-\mu_{R}\right)}{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)+\theta_{A}\left(p,k\right)}\right)=-\operatorname{E}\left(\frac{\theta_{A}\left(p,k\right)\theta_{R}\left(k,p\right)-\theta_{A}\left(p,k\right)\mu_{R}}{\sum_{\left(q,w\right)\in P}\theta_{A}\left(q,w\right)+\theta_{A}\left(p,k\right)}\right)$
Again applying that $E\left(\frac{X}{Y}\right)\approx\frac{E(X)}{E(Y)}$, the expression can be simplified to
$D_{\mu}=-\frac{\mu_{A}\left(p\right)}{\left(N-1\right)*\mu_{A}+\mu_{A}\left(p\right)}\left(\mu_{R}\left(p\right)-\mu_{R}\right)$
Further invoking the approximation of equation 4,
$D_{\mu}=-\frac{\mu_{A}\left(p\right)}{N\mu_{A}}\left(\mu_{R}\left(p\right)-\mu_{R}\right)$
A.3.2 Variance of the differential
The variance requires the Taylor expansion, which is [8]
$\operatorname{Var}\left(\frac{X}{Y}\right)\approx\frac{\operatorname{Var}\left(X\right)}{\operatorname{E}\left(Y\right)^{2}}-2*\frac{\operatorname{Cov}\left(X,Y\right)*\operatorname{E}\left(X\right)}{E\left(Y\right)^{3}}+\frac{\operatorname{E}\left(X\right)^{2}*\operatorname{Var}\left(Y\right)}{\operatorname{E}\left(Y\right)^{4}}$
In the free throw case for team $B$, this becomes
$O_{\sigma_{B}}^{2}=\frac{\operatorname{Var}\left(\sum_{q\in T_{B}}\mu_{A}\left(q\right)\mu_{R}\left(q\right)\right)}{\left(N\mu_{A}\right)^{2}}$ $-2\frac{\operatorname{Cov}\left(\sum_{q\in T_{B}}\mu_{A}\left(q\right)\mu_{R}\left(q\right),\sum_{q\in T_{B}}\mu_{A}\left(q\right)\right)*N\mu_{A}\mu_{R}}{\left(N\mu_{A}\right)^{3}}$
$+\frac{\left(N\mu_{R}\mu_{A}\right)^{2}*\operatorname{Var}\left(\sum_{q\in T_{B}}\mu_{A}\left(q\right)\right)}{\left(N\mu_{A}\right)^{4}}$
Or
$O_{\sigma_{B}}^{2}=\frac{\operatorname{Var}\left(\sum_{q\in T_{B}}\mu_{A}\left(q\right)\mu_{R}\left(q\right)\right)}{\left(N\mu_{A}\right)^{2}}$ $-2\frac{\operatorname{Cov}\left(\sum_{q\in T_{B}}\mu_{A}\left(q\right)\mu_{R}\left(q\right),\sum_{q\in T_{B}}\mu_{A}\left(q\right)\right)\mu_{R}}{\left(N\mu_{A}\right)^{2}}$
$+\frac{\mu_{R}^{2}*\operatorname{Var}\left(\sum_{q\in T_{B}}\mu_{A}\left(q\right)\right)}{\left(N\mu_{A}\right)^{2}}$
Some basic facts about variance can be applied to the expression. $C\operatorname{Var}\left(X\right)=\operatorname{Var}\left(\sqrt{C}X\right)$, and $C\operatorname{Cov}\left(X,Y\right)=\operatorname{Cov}\left(\sqrt{C}X,\sqrt{C}Y\right)=\operatorname{Cov}\left(X,CY\right)$, so the equation can be simplified to
$O_{\sigma_{B}}^{2}=$ $\frac{\operatorname{Var}\left(\sum_{q\in T_{B}}\frac{\mu_{A}\left(q\right)}{\mu_{A}}\mu_{R}\left(q\right)\right)}{N^{2}}-\frac{2\operatorname{Cov}\left(\sum_{q\in T_{B}}\frac{\mu_{A}\left(q\right)}{\mu_{A}}\mu_{R}\left(q\right),\sum_{q\in T_{B}}\frac{\mu_{A}\left(q\right)}{\mu_{A}}\mu_{R}\right)}{N^{2}}$
$+\frac{\operatorname{Var}\left(\sum_{q\in T_{B}}\frac{\mu_{A}\left(q\right)}{\mu_{A}}\mu_{R}\right)}{N^{2}}$
It is known that $\operatorname{Var}\left(X\right)+\operatorname{Var}\left(Y\right)-2\operatorname{Cov}\left(X,Y\right)=\operatorname{Var}\left(X-Y\right)$ so this can be rewritten to
$O_{\sigma_{B}}^{2}=\frac{\operatorname{Var}\left(\sum_{q\in T_{B}}\frac{\mu_{A}\left(q\right)}{\mu_{A}}*\mu_{R}\left(q\right)-\frac{\mu_{A}\left(q\right)}{\mu_{A}}*\mu_{R}\right)}{N^{2}}$
Or
$O_{\sigma_{B}}^{2}=\frac{\operatorname{Var}\left(\sum_{q\in T_{B}}\frac{\mu_{A}\left(q\right)}{\mu_{A}}*\left(\mu_{R}\left(q\right)-\mu_{R}\right)\right)}{N^{2}}=\frac{\sum_{q\in T_{B}}\operatorname{Var}\left(\frac{\mu_{A}\left(q\right)}{\mu_{A}}*\left(\mu_{R}\left(q\right)-\mu_{R}\right)\right)}{N^{2}}$
There are $N$ players that contribute identical variance, so this can be rewritten to
$O_{\sigma_{B}}^{2}=\frac{N*\operatorname{Var}_{q\in Q,w\in W}\left(\frac{\theta_{A}\left(q,w\right)}{\mu_{A}}*\left(\theta_{R}\left(q,w\right)-\mu_{R}\right)\right)}{N^{2}}$
For the time being, define
$\sigma_{q}^{2}=\operatorname{Var}_{q\in Q,w\in W}\left(\frac{\theta_{A}\left(q,w\right)}{\mu_{A}}*\left(\theta_{R}\left(q,w\right)-\mu_{R}\right)\right)$
$\sigma_{p}^{2}=\operatorname{Var}_{w\in W}\left(\frac{\theta_{A}\left(p,w\right)}{\mu_{A}}*\left(\theta_{R}\left(p,w\right)-\mu_{R}\right)\right)$
The variance for team $B$ is then
$O_{\sigma_{B}}^{2}=\frac{N}{N^{2}}\sigma_{q}^{2}=\frac{\sigma_{q}^{2}}{N}$
For team $A$, the variance is
$O_{\sigma_{A}}^{2}=$ $\frac{\operatorname{Var}\left(\sum_{q\in T_{A}}\mu_{A}\left(q\right)\mu_{R}\left(q\right)\right)}{\left(\left(N-1\right)\mu_{A}\right)^{2}}$
$-2\frac{\operatorname{Cov}\left(\sum_{q\in T_{A}}\mu_{A}\left(q\right)\mu_{R}\left(q\right),\sum_{q\in T_{A}}\mu_{A}\left(q\right)\right)*\left(\left(N-1\right)\mu_{A}\mu_{R}+\mu_{A}\left(p\right)\mu_{R}\left(p\right)\right)}{\left(N\mu_{A}\right)^{3}}$
$+\frac{\left(\left(N-1\right)\mu_{A}\mu_{R}+\mu_{A}\left(p\right)\mu_{R}\left(p\right)\right)^{2}*\operatorname{Var}\left(\sum_{q\in T_{A}}\mu_{A}\left(q\right)\right)}{\left(N\mu_{A}\right)^{4}}$
Because of the assumption of equation 5 that $\left(N-1\right)\mu_{A}\mu_{R}+\mu_{A}\left(p\right)\mu_{R}\left(p\right)\approx N\mu_{R}\mu_{A}$, this becomes
$O_{\sigma_{A}}^{2}=$ $\frac{\operatorname{Var}\left(\sum_{q\in T_{A}}\mu_{A}\left(q\right)\mu_{R}\left(q\right)\right)}{\left(N\mu_{A}\right)^{2}}-2\frac{\operatorname{Cov}\left(\sum_{q\in T_{A}}\mu_{A}\left(q\right)\mu_{R}\left(q\right),\sum_{q\in T_{A}}\mu_{A}\left(q\right)\right)*\mu_{R}}{\left(N\mu_{A}\right)^{2}}$
$+\frac{\left(N\mu_{A}\mu_{R}\right)^{2}*\operatorname{Var}\left(\sum_{q\in T_{A}}\mu_{A}\left(q\right)\right)}{\left(N\mu_{A}\right)^{4}}$
Again some basic facts about variance can be applied to the expression. $C\operatorname{Var}\left(X\right)=\operatorname{Var}\left(\sqrt{C}X\right)$, and $C\operatorname{Cov}\left(X,Y\right)=\operatorname{Cov}\left(\sqrt{C}X,\sqrt{C}Y\right)=\operatorname{Cov}\left(X,CY\right)$, so the equation can be simplified to
$O_{\sigma_{A}}^{2}=$ $\frac{\operatorname{Var}\left(\sum_{q\in T_{A}}\frac{\mu_{A}\left(q\right)}{\mu_{A}}\mu_{R}\left(q\right)\right)}{N^{2}}$
$-\frac{2\operatorname{Cov}\left(\sum_{q\in T_{A}}\frac{\mu_{A}\left(q\right)}{\mu_{A}}\mu_{R}\left(q\right),\sum_{q\in T_{A}}\frac{\mu_{A}\left(q\right)}{\mu_{A}}\mu_{R}\right)}{N^{2}}+\frac{\operatorname{Var}\left(\sum_{q\in T_{A}}\frac{\mu_{A}\left(q\right)}{\mu_{A}}\mu_{R}\right)}{N^{2}}$
It is known that $\operatorname{Var}\left(X-Y\right)=\operatorname{Var}\left(X\right)+\operatorname{Var}\left(Y\right)-2\operatorname{Cov}\left(X,Y\right)$ so this can be rewritten to
$O_{\sigma_{A}}^{2}=\frac{\operatorname{Var}\left(\sum_{q\in T_{A}}\frac{\mu_{A}\left(q\right)}{\mu_{A}}*\mu_{R}\left(q\right)-\frac{\mu_{A}\left(q\right)}{\mu_{A}}*\mu_{R}\right)}{N^{2}}$
In this case, there are $\left(N-1\right)$ players that contribute variance $\sigma_{q}^{2}$ and one that contributes $\sigma_{p}^{2}$, so this becomes
$O_{\sigma_{A}}^{2}=\frac{\left(N-1\right)*\sigma_{q}^{2}+\sigma_{p}^{2}}{N^{2}}$
The variance of the differential is the variance of team A’s score plus the variance of team B’s score.
$D_{\sigma}^{2}=\frac{\left(N-1\right)*\sigma_{q}^{2}+\sigma_{p}^{2}}{N^{2}}+\frac{\sigma_{q}^{2}}{N}=\frac{\left(2N-1\right)\sigma_{q}^{2}+\sigma_{p}^{2}}{N^{2}}$
Now $\sigma_{q}^{2}$ and $\sigma_{p}^{2}$ must be calculated. Starting with $\sigma_{q}^{2}$ , it is
$\sigma_{q}^{2}=\operatorname{Var}_{q\in Q,w\in W}\left(\frac{\theta_{A}\left(q,w\right)}{\mu_{A}}\left(\theta_{R}\left(q,w\right)-\mu_{R}\right)\right)$
$\sigma_{q}^{2}$ in this form can be easily computed if $W$ is available as a discrete set. It can also be translated to a more understandable form.
Since $\operatorname{Var}\left(X\right)=\operatorname{E}\left(X^{2}\right)+\operatorname{E}\left(X\right)^{2}$ and the expected value of $\frac{\theta_{A}\left(q,w\right)}{\mu_{A}}\left(\theta_{R}\left(q,w\right)-\mu_{R}\right)=0$, this becomes
$\sigma_{q}^{2}=\operatorname{E}_{q\in Q,w\in W}\left(\frac{\theta_{A}\left(q,w\right)^{2}}{\mu_{A}^{2}}\left(\theta_{R}\left(q,w\right)-\mu_{R}\right)^{2}\right)$
$=\operatorname{E}_{q}\left(\operatorname{E}_{w}\left(\frac{\theta_{A}\left(q,w\right)^{2}}{\mu_{A}^{2}}\left(\theta_{R}\left(q,w\right)-\mu_{R}\right)^{2}\right)\right)$
$\operatorname{E}\left(X^{2}\right)=\operatorname{Var}\left(X\right)-\operatorname{E}\left(X\right)^{2}$ so this can be translated into
$\sigma_{q}^{2}=\operatorname{E}_{q}\left(\operatorname{Var}_{w}\left(\frac{\theta_{A}\left(q,w\right)}{\mu_{A}}\left(\theta_{R}\left(q,w\right)-\mu_{R}\right)\right)+\operatorname{E}_{w}\left(\frac{\theta_{A}\left(q,w\right)}{\mu_{A}}\left(\theta_{R}\left(q,w\right)-\mu_{R}\right)\right)^{2}\right)$
$=\ \operatorname{E}_{q}\left(\operatorname{Var}_{w}\left(\frac{\theta_{A}\left(q,w\right)}{\mu_{A}}\left(\theta_{R}\left(q,w\right)-\mu_{R}\right)\right)+\operatorname{E}_{w}\left(\frac{\theta_{A}\left(q,w\right)\theta_{R}\left(q,w\right)}{\mu_{A}}-\frac{\theta_{A}\left(q,w\right)\mu_{R}}{\mu_{A}}\right)^{2}\right)$
$=\operatorname{E}_{q}\left(\operatorname{Var}_{w}\left(\frac{\theta_{A}\left(q,w\right)}{\mu_{A}}\left(\theta_{R}\left(q,w\right)-\mu_{R}\right)\right)+\left(\operatorname{E}_{w}\left(\frac{\theta_{A}\left(q,w\right)\theta_{R}\left(q,w\right)}{\mu_{A}}\right)-\operatorname{E}_{w}\left(\frac{\theta_{A}\left(q,w\right)\mu_{R}}{\mu_{A}}\right)\right)^{2}\right)$
$=\operatorname{E}_{q}\left(\operatorname{Var}_{w}\left(\frac{\theta_{A}\left(q,w\right)}{\mu_{A}}\left(\theta_{R}\left(q,w\right)-\mu_{R}\right)\right)+\left(\frac{\mu_{A}\left(q\right)\mu_{R}\left(q\right)}{\mu_{A}}-\frac{\mu_{A}\left(q\right)\mu_{R}}{\mu_{A}}\right)^{2}\right)$
$=\operatorname{E}_{q}\left(\operatorname{Var}_{w}\left(\frac{\theta_{A}\left(q,w\right)}{\mu_{A}}\left(\theta_{R}\left(q,w\right)-\mu_{R}\right)\right)+\left(\frac{\mu_{A}\left(q\right)}{\mu_{A}}\left(\mu_{R}\left(q\right)-\mu_{R}\right)\right)^{2}\right)$
Separating out the two terms yields
$\sigma_{q}^{2}=$ $\left(\operatorname{E}_{q}\operatorname{Var}_{w}\left(\frac{\theta_{A}\left(q,w\right)}{\mu_{A}}\left(\theta_{R}\left(q,w\right)-\mu_{R}\right)\right)+\operatorname{E}_{q}\left(\frac{\mu_{A}\left(q\right)}{\mu_{A}}\left(\mu_{R}\left(q\right)-\mu_{R}\right)\right)^{2}\right)$
$=\operatorname{E}_{q}\left(\operatorname{Var}_{w}\left(\frac{\theta_{A}\left(q,w\right)}{\mu_{A}}\left(\theta_{R}\left(q,w\right)-\mu_{R}\right)\right)\right){N}+\operatorname{Var}_{q}\left(\frac{\mu_{A}\left(q\right)}{\mu_{A}}\left(\mu_{R}\left(q\right)-\mu_{R}\right)\right)$
$\quad+\left(\operatorname{E}\left(\frac{\mu_{A}\left(q\right)}{\mu_{A}}\left(\mu_{R}\left(q\right)-\mu_{R}\right)\right)\right)^{2}$
Note that $\operatorname{E}\left(\frac{\mu_{A}\left(q\right)}{\mu_{A}}\left(\mu_{R}\left(q\right)-\mu_{R}\right)\right)=\operatorname{E}\left(\frac{\mu_{A}\left(q\right)*\mu_{R}\left(q\right)}{\mu_{A}}\right)-\operatorname{E}\left(\frac{\mu_{A}\left(q\right)*\mu_{R}}{\mu_{A}}\right)=\mu_{R}-\frac{\mu_{A}*\mu_{R}}{\mu_{A}}=\mu_{R}-\mu_{R}=0$. That leads to a form of
$\sigma_{q}^{2}=\left(\operatorname{E}_{q}\left(\operatorname{Var}_{w}\left(\frac{\theta_{A}\left(q,w\right)}{\mu_{A}}\left(\theta_{R}\left(q,w\right)-\mu_{R}\right)\right)\right)+\operatorname{Var}_{q}\left(\left(\frac{\mu_{A}\left(q\right)}{\mu_{A}}\left(\mu_{R}\left(q\right)-\mu_{R}\right)\right)\right)\right)$
The first term matches $\tau_{R}^{2}$ and the second term matches $\sigma_{R}^{2}$. The resulting overall variance is then
$\sigma_{q}^{2}=\sigma_{R}^{2}+\tau_{R}^{2}$
For $\sigma_{p}^{2}$, because of equation 2
$\sigma_{p}^{2}=\tau_{R}\left(p\right)^{2}\approx\tau_{R}^{2}$
And finally
$D_{\sigma}^{2}==\frac{\left(2N-1\right)\sigma_{q}^{2}+\sigma_{p}^{2}}{N^{2}}=\frac{\left(2N-1\right)\left(\sigma_{R}^{2}+\tau_{R}^{2}\right)+\tau_{R}^{2}}{N^{2}}=\frac{\left(2N-1\right)\sigma_{R}^{2}+2N\tau_{R}^{2}}{N^{2}}$
A.3.3 Differential distribution
Now the normal approximation CDF at zero can be written
$\frac{1}{2}\left[1+\operatorname{erf}\left(\frac{\frac{\mu_{A}\left(p\right)}{N\mu_{A}}\left(\mu_{R}\left(p\right)-\mu_{R}\right)}{\sqrt{2*\frac{\left(2N-1\right)\sigma_{R}^{2}+2N\tau_{R}^{2}}{N^{2}}}}\right)\right]$
Applying the Taylor first-order approximation yields
$\frac{1}{2}\left[1+\frac{2}{\sqrt{\pi}}*\frac{\frac{\mu_{A}\left(p\right)}{N\mu_{A}}\left(\mu_{R}\left(p\right)-\mu_{R}\right)}{\sqrt{2*\frac{\left(2N-1\right)\sigma_{R}^{2}+2N\tau_{R}^{2}}{N^{2}}}}\right]$
$=\frac{1}{2}\left[1+\frac{2}{\sqrt{\pi}}*\frac{\frac{\mu_{A}\left(p\right)}{\mu_{A}N}\left(\mu_{R}\left(p\right)-\mu_{R}\right)}{\sqrt{2*\frac{\left(2N-1\right)\sigma_{R}^{2}+2N\tau_{R}^{2}}{N^{2}}}}\right]$
$=\frac{1}{2}\left[1+\frac{1}{\sqrt{\pi\left(N-\frac{1}{2}\right)}}*\frac{\frac{\mu_{A}\left(p\right)}{\mu_{A}}\left(\mu_{R}\left(p\right)-\mu_{R}\right)}{\sqrt{\sigma_{R}^{2}+\frac{2N}{2N-1}\tau_{R}^{2}}}\right]$
A.4 Total across categories
Adding together the G-scores for all counting statistics percentage statistics yields
$\frac{1}{2}\left[|C|+\frac{1}{\sqrt{\pi*\left(N-\frac{1}{2}\right)}}*\sum_{c}G_{c}\right]$
Where $G_{c}$ is the G-score for a category.
This is player value, as defined in Section 2.1. G-score is clearly proportional to it, with each point of G-score being worth
$\frac{1}{2\sqrt{\pi*\left(N-\frac{1}{2}\right)}}$
So G-score meets the criteria for player value as well
Appendix B Position requirement
Most leagues have position requirements, which do not factor into the justification for G-scores. For fantasy basketball, the omission is not enormously problematic because there are often many flex spots, players are eligible for many positions, and value tends to be spread out fairly evenly between positions anyway. See Table 8 for the distribution from the 2022-2023 season.
Center Point Guard Shooting Guard Power Forward Small Forward
Rank
1-52 13 12 10 9 8
53-104 10 11 13 11 6
105-156 7 10 11 15 8
Overall 30 34 34 36 22
Table 8: Top 156 players by G-score in 2023
If a league has particularly tight position requirements, then a value-over-replacement (VOR) approach across positions may be warranted. VOR adjusts values by estimating how valuable the best player available after the draft would be for each position, then subtracts those values from all players of those positions. It is commonly used for fantasy football, which has tight position requirement [9]
Appendix C Change in performance distributions over time
Performance distributions can change throughout a season. In lieu of a comprehensive model to account for any kind of performance drift rigorously, two simplified models are presented here to account for two major kinds of change over time- injury and random performance drift
C.1 Injuries
Sometimes, players are too injured to play for extended periods of time. As unfortunate as this is, it is a part of fantasy basketball.
The obvious way to model injury risk is to include the probability of injury into the probability distributions of player performances represented by $W$. This is not entirely satisfactory because leagues often have designed spots for injured players, allowing managers to replace them with replacement-level alternatives under certain conditions. An ideal treatment of injury risk would account for that, and ameliorate the expected effect of injury accordingly. Additionally, if a manager is concerned more with winning a league than with placing well in general, they may want to assume good health of their players, because their chances of winning would be minimal otherwise.
These complications motivate a treatment of injury risk which is parameterized and adjustable based on circumstances. One solution is scaling injury risk down by a parameter, then calculating scores with replacement value added in and scaled down by a second parameter.
In math, first player statistics are modified
$I^{\prime}=I*\upsilon$
$\mu_{c}^{\prime}=\mu_{c}*(1-I^{\prime})$
Where $\upsilon$ is a factor between zero and one, $I^{\prime}$ is an adjusted injury rate $I$, and $\mu_{c}^{\prime}$ is an adjusted mean of a counting statistic $\mu_{c}$. In this context volume statistics such as free throws attempted are included as counting statistics, since they need to be adjusted as well. Note that this adjustment should not be included in the calculation of coefficients for scoring, since it does not affect aggregate injury expectation.
Then, after scores are calculated, they are adjusted based on replacement value
$X^{\prime}=X+I^{\prime}*\psi*R$
Where $\psi$ is the expected fraction of games for which an injured player can be replaced, $R$ is replacement value, and $X^{\prime}$ is a modified score $X$.
Replacement-level value can be approximated in several ways. One is to spread value evenly, except with a negative for turnovers. So for example if the replacement value is $-3$, by category it is broken down by $R=-\frac{3}{7}$ for non-turnover categories and $R=\frac{3}{7}$ for turnovers. The sum is $-3$ as needed
C.2 Performance Drift
Player performances inevitably drift over time. When attentive managers are allowed to replace under-performing players, they can replace low-level players that dip below replacement-level value. This dilutes the value of low-level players who are likely to be replaced, in a way that can be analyzed mathematically
C.2.1 A simple model of performance drift
Assume that the replacement value remains the same over the course of a season. This is intuitively reasonable; individual players may go up or down but the value of the player at rank e.g. 157 should be relatively consistent.
Further, assume that each player’s value is perturbed by noise $\epsilon$ with mean zero and standard deviation $\sigma$ by some point during the season. If $\epsilon$ drops the player’s value below the replacement value, then they are not worth having on a team and will be dropped in favor of a player with exactly the replacement value.
Consider the expected performance value of a player above replacement to be $\mu$, with perturbation $\epsilon_{f}$. Then, their perturbed value $F$ is
$F=\begin{cases}\mu+\epsilon_{f},&\text{if }\mu+\epsilon_{f}\geq 0\\ 0,&\text{otherwise}\end{cases}$
Without any investment into a particular roster slot, a manager can always choose a replacement-level player for essentially no cost. Players with exactly replacement value before perturbation ($\mu=0$) have post-perturbation value
$S=\begin{cases}\epsilon_{s},&\text{if }\epsilon_{s}\geq 0\\ 0,&\text{otherwise}\end{cases}$
Since picking a player removes a spot that could be used for a costless replacement-level player or “streamer”, the real value of picking a player is $E[F-S]=E[F]-E[S]$
C.2.2 SAVOR
It is posited that a novel metric called SAVOR is equivalent to $E[F]-E[S]$ given the assumptions. SAVOR-adjusted player value given starting point $\mu$ is
$\mu*\phi(\frac{\mu}{\sigma})-\frac{\sigma}{\sqrt{2\pi}}\left(1-e^{\frac{-\mu^{2}}{2\sigma^{2}}}\right)$
Where $\phi$ is the cumulative distribution function of a standard Normal.
This adjustment is monotonic with respect to $\mu$, so it only has relevance to auctions, where value is important in addition to order. Table 9 shows an example of its application. As expected, the adjustment dilutes the value of low-level players, to a degree that increases with $\sigma$
Pre-Savor Post-SAVOR
$\mu$ $\sigma=0$ $\sigma=0.1$ $\sigma=1$
Value Percent Value Percent Value Percent Value Percent
0 0% 0 0% 0 0% 0 0%
1 10% 1 10% 0.96 9.76% 0.68 8.05%
2 20% 2 20% 1.96 19.92% 1.61 18.94%
3 30% 3 30% 2.96 30.08% 2.60 30.62%
4 40% 4 40% 3.96 40.24% 3.60 42.38%
Table 9: Results of a SAVOR calculation, re-normalized to sum to 100
C.2.3 Justifying SAVOR
With $\epsilon_{f}$ and $\epsilon_{s}$ defined as normal errors with standard deviation $\sigma$,
$F=\begin{cases}\mu+\epsilon_{f},&\text{if }\mu+\epsilon_{f}\geq 0\\ 0,&\text{otherwise}\end{cases}$
$S=\begin{cases}\epsilon_{s},&\text{if }\epsilon_{s}\geq 0\\ 0,&\text{otherwise}\end{cases}$
Using the definition of the expected value and the probability density of the normal distribution,
[…]
Changing the variable in the expression for $E[F]$ from $x$ to $y+\mu$ yields
[…]
Therefore
[…]
The second part of the expression can be identified as $\mu*\phi(\frac{\mu}{\sigma})$. This simplifies to
[…]
Substituting $-\sigma^{2}*d(-\frac{y^{2}}{2\sigma^{2}})$ for $ydy$ yields
$E[F]-E[S]=\mu*\phi(\frac{\mu}{\sigma})-\int_{-\mu}^{0}\frac{\sigma}{\sqrt{2\pi}}e^{\frac{-y^{2}}{2\sigma^{2}}}d(-\frac{y^{2}}{2\sigma^{2}})$
Upon integration, this yields
$E[F]-E[S]=\mu*\phi(\frac{\mu}{\sigma})-\frac{\sigma}{\sqrt{2\pi}}\left(e^{\frac{-y^{2}}{2\sigma^{2}}}|_{-\mu}^{0}\right)=\mu*\phi(\frac{\mu}{\sigma})-\frac{\sigma}{\sqrt{2\pi}}\left(1-e^{\frac{-\mu^{2}}{2\sigma^{2}}}\right)$